OpenStudy (stormoshawty):

Choose 1 set of points from the choices below. Then, solve the problem and post your solution, showing your steps. Write an equation in point-slope form for the line that passes through one of the following pairs of points (you may choose the pair you want to work with). Then, use the same set of points to write the equation in standard form and again in slope-intercept form. Point pairs (5, 1), (–3, 4) (0, –2), (3, 2) (–2, –1), (1, 2)

OpenStudy (shamil98):

Find the slope first. \[\large \frac{ y_2-y_1 }{ x_2-x_1 }\]

OpenStudy (stormoshawty):

I used the first one so the slope is \[\frac{ 1 }{ 4 }\] i'm pretty sure unless I did it wrong.

OpenStudy (shamil98):

4-1/-3-5 = 3/-8

OpenStudy (stormoshawty):

Oh. Yeah, I put 5 on top on accident.

OpenStudy (shamil98):

Okay. Now, (5, 1) and -3/8 input those into \[\large y-y_1=m(x-x_1)\]

OpenStudy (shamil98):

\[\large y - 1 = -\frac{ 3 }{ 8 } (x-5)\]

OpenStudy (shamil98):

and there you go.

OpenStudy (stormoshawty):

I got that. now solve that?

OpenStudy (shamil98):

Yeah, change it to slope-intercept.

OpenStudy (shamil98):

Then from there change it to standard form.

OpenStudy (stormoshawty):

Oh my god. So much work!!!!!!!!!

OpenStudy (shamil98):

yep

OpenStudy (stormoshawty):

it says standard form then slope intercept form.

OpenStudy (shamil98):

You can do either way.. but ok \[\huge y-1= \frac{ -3(x-5) }{ 8 } \] \[\huge y - 1 = \frac{ -3x+15 }{ 8 }\] \[\huge 8y-8 = -3x + 15\] \[\huge 3x + 8y = 23\]

OpenStudy (shamil98):

From there just change it to slope-intercept.

OpenStudy (stormoshawty):

Is that Regular form?

OpenStudy (shamil98):

Yes, that is standard form..

OpenStudy (stormoshawty):

Standard. Whatever. Lol. That's what i meant.